alones (alones):

Evaluate /:

alones (alones):

−3√-8

alones (alones):

I got −1√-2 for my answer

OpenStudy (bobo-i-bo):

Do you know what \[\sqrt{-1} \] is?

alones (alones):

Umm. no

alones (alones):

it's squred root

OpenStudy (bobo-i-bo):

\[\sqrt{-1}=i\] where i is THE imaginary number... have you covered complex numbers at all?

rebeccaxhawaii (rebeccaxhawaii):

mm

alones (alones):

ofc i did.

OpenStudy (bobo-i-bo):

Okay... then notice that \[-3\sqrt{-8}=-3\sqrt{8}\sqrt{-1}\]

Directrix (directrix):

@Bobo-i-bo Is that ^^^ the final answer?

OpenStudy (bobo-i-bo):

No, i wouldn't give you the final answer :P

OpenStudy (bobo-i-bo):

But i'm trying to help you, lol

alones (alones):

I said my final answer already Hm._.

OpenStudy (bobo-i-bo):

This will help you out: http://www.mathwords.com/s/square_root_rules.htm

alones (alones):

btw i'd like a person to explain me not this sites /;

alones (alones):

@Bobo-i-bo so that was your final answer -6i√2

OpenStudy (bobo-i-bo):

yes

alones (alones):

welp that doesn't not really looks liek the right asnwer :#

Directrix (directrix):

What about the explanation now that the assumed correct answer is known. @Bobo-i-bo

alones (alones):

#rekt

OpenStudy (bobo-i-bo):

\[-3\sqrt{-8}=-3\sqrt{8}\sqrt{-1}=-3\sqrt{4}\sqrt{2}i=-3*2\sqrt{2}i=-6\sqrt{2}i\]

OpenStudy (bobo-i-bo):

Understand it all?

OpenStudy (bobo-i-bo):

It follows from the property that, given numbers a and b, \[\sqrt{ab}=\sqrt{a}\sqrt{b}\]

alones (alones):

well i got the difrent answer /:

OpenStudy (bobo-i-bo):

and how did you get your answer? please may i see working?

alones (alones):

I cannot solve this since it's undetifined